OpenStudy (anonymous):

What is the volume of 700.0 g of ether if the density of ether is 0.70 g/mL? How would I solve this? I always get confused with problems like this. Please help and explain!

OpenStudy (cwrw238):

you just need to remember the relation between mass, volume and density density = mass/volume

OpenStudy (anonymous):

So would I go 0.70 g/mL = 700.0g/volume mL ?

OpenStudy (cwrw238):

yes 0.70 = 700 / V where V is in mL

OpenStudy (anonymous):

Then do the mL cancel out or do I just divide the density with the mass?

OpenStudy (cwrw238):

cross multiply:- 0.70V = 700 divide both sides by 0.70 V = 700 / 0.70

OpenStudy (anonymous):

so it would be 1.000?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

So that would be the volume of ether in the problem?

OpenStudy (cwrw238):

yes 1,000 mL

OpenStudy (anonymous):

I got it. I understand it better thank you so much for helping me!

OpenStudy (cwrw238):

ok